Find an equation that passes through (3,5) that is perpendicular to the line 3x+6y=4
equation that passes passes through (3,5 that perpendicular to the line 3x+6y=4 is
Answer #1

Ani find the equation of the line: 3x+6y=4 GYE 4-32 6 y = -39+4. y - 30 tu y: an + 62 6s 6 42 y = -2 / + dlo n 3 is in slope